OpenStudy (anonymous):

The wavelength of a light wave will affect the light’s Frequency Amplitude Pitch Color

OpenStudy (calculator):

\[c= f \lambda\]

OpenStudy (calculator):

\[i \hbar \frac{\partial \Psi}{\partial t}= - \frac{\hbar^2}{2m} \frac{\partial ^2 \Psi}{\partial x^2} \\ \\ \Psi(x,t)= Ae^{i(kx-\omega t)} \\ \\ i \hbar \frac{\partial \Psi}{\partial t}= i \hbar (-i \omega)Ae^{i(kx-\omega t)}= \hbar \omega Ae^{i(kx-\omega t)} \\ \\ \text{gives} \\ \\ - \frac{\hbar^2}{2m} \frac{\partial ^2 \Psi}{\partial x^2} = \frac{\hbar^2k^2}{2m}Ae^{i(kx-\omega t)}\]

OpenStudy (anonymous):

The wavelength of light affects the frequency. The frequency affects the color, so it's sort of a trick question. @Calculator what does all that mean? and what physics class is it from?

OpenStudy (anonymous):

This is such an ambiguous question... It's like asking "The wavelength of a soundwave will affect the sound's: Frequency - Amplitude - Pitch - Color" Wavelength is a function of frequency so I wouldn't call it an affectation per se. In the case of light, I'd say the wavelength affects color. In the case of sound, it would affect pitch.

OpenStudy (anonymous):

if affects color and it never affects the frequency or the pitch... the frequency solely depends on the source.. once the source produces a wave of a particular frequency it ll never change (ofcourse not considering damping)

OpenStudy (aaronq):

^ what if the light is colourless to the human eye ;)

OpenStudy (anonymous):

you mean outside our visible range ? :P.. still i holds good.. when we say color.. we mean wavelength :P

OpenStudy (aaronq):

nope! that is an unfair generalization what colour is the light with wavelength of 1 m, huh? rubbish! lol

OpenStudy (shane_b):

The *best* answer given the wording would be color. I concur with @ qweqwe123123123123111 's comment.

OpenStudy (anonymous):

@aarong our eyes are not sensitive to such wavelengths.. and thus we cannot see them and hence we can't assign a color to them :P but i think that lies in the microwave region!..

OpenStudy (aaronq):

so you're saying that were unable to characterize all wavelengths or frequencies of light by colour because our eyes are not sensitive.. so the correct answer would be frequency.

OpenStudy (anonymous):

@ Peter14 What he wrote in his second post is the Schrödinger equation of a particle in a free field, which is not related to the question. Physics courses in university offer quantum mechanics which will cover this equation.

OpenStudy (anonymous):

@aaronq what are you even talking about..? :P.. the question is what does the wavelength of the light affect.. and thats its color.. frequency is a constant

OpenStudy (aaronq):

we can say that the wavelength of light affects it's colour when we can qualitatively characterize it, since we can't do that outside of the visible range it is illogical to use "colours" to describe it. frequency is not a constant... it's inversely proportional to wavelength, you know this, come on. \[\nu=\frac{ c }{ \lambda } \]

OpenStudy (anonymous):

frequency is a constant why is everyone mistaken by the v= f(lambda) equation?? f is a constant.. just like how h is a constant in E = hf you can say h = E/f and is dependent on E and f?? but f is not a universal constant.. its a constant that depends on the source... Once the source emits a wave of particular frequency it ll never change regardless of which medium it is travelling in ! and we HAVE quantitatively characterized it by USING WAVELENGTH itself.. but m not so sure anymore about my answer cause wavelength doesn't really decide the color, its the frequency.. i had raised a question on it :P.. so i dunno now :P

OpenStudy (anonymous):

Wavelength affects frequency. Color is a construction of the human eye.

OpenStudy (anonymous):

Wavelength does not affect frequency wavelength affects the velocity of the wave!!

OpenStudy (anonymous):

but the velocity of the wave is constant

OpenStudy (anonymous):

Hogwash. Wavelength affects frequency like the number 10 affects the fraction 1/10th. Wavelength and frequency are the same damned thing looked at from 2 different directions. Also, if frequency were a "constant", then tell me why it's represented as a VARIABLE using lambda! And don't confuse the words AFFECT and EFFECT, either. You EFFECTively change wavelength by changing frequency and vice-versa, but what that change AFFECTS is COLOR in the visible segment of the electromagnetic spectrum.

OpenStudy (anonymous):

huh. we're still arguing over this and the original question was asked 2 days ago and probably has been turned in.

OpenStudy (anonymous):

Heheh...true, that. :-) And I'd give my left arm to know what the prof deemed the "right" answer... :-)

OpenStudy (anonymous):

I wouldn't go that far, I'd just give a tooth or two. More replaceable.

OpenStudy (anonymous):

I could use the technology the brits are developing. They recently came up with a way of growing new teeth in mice. The americans came up with a crude system awhile back that can regrow dentin but not the enamel.

OpenStudy (anonymous):

you can't change wavelength by changing frequency!! .. you can't even change frequency by changing wavelength!!..

OpenStudy (anonymous):

and wavelength change doesn't AFFECT color :P.. its the frequency that matters you can have a radiation having different wavelengths but same frequency!!.. and they would cause same color cause color is basically how many waves pass your photoreceptors per second.. i just had discussion about that.. !

OpenStudy (anonymous):

-"you can't change wavelength by changing frequency!! .. you can't even change frequency by changing wavelength!!" [facepalm] This is a joke, yes?

OpenStudy (anonymous):

Nope not a joke.. pure fact tell me this.. if you wanna say 'a' depends on 'b'.. then as 'b' changes 'a' should change correct?? so give me one case where you change wavelength and frequency changes and then you can say freq depends on wavelength i ll give you refraction case where wavelenght changes but nothing happens to frequency.. go beyond equations.. why do you think raman scattering and compton scattering were a nightmare to the classical wave theory?? why can't it explain INCOHERENT scattering?.. cause the scattered light had a change in frequency.. why? that was such mystery?? if frequency was dependent on wavelength you can easily explain those using simple wave theory... so if you wanna know what frequency actually depends on , its the energy!! electromagnetic waves are created by oscillating charges, and have the same frequency as oscillating charges. what decides the frequency of the oscillating charges ? depends on what energy you put into them.. and thats were the whole quantization and the quantum mechanics came into the picture.. but its ENERGY of the oscillating charges that decide the FREQUENCY of the radiation that comes out.. and once the frequency is set it will not change! unless it starts having microscopic interactions with electrons and loses or gains energy (thats how the incoherent scatterings are explained by photon theory of light)!